# Adsorption of volatiles on dust grains in protoplanetary disks

### v2

@@ 1 Introduction
The adsorption of volatile chemical species onto the surfaces of dust grains is of broad and critical interest to researchers studying the interstellar medium (ISM), star formation, and planet formation. For example, in protoplanetary disks (PPDs hereafter), dusts may play a central role in controlling the thermochemical conditions. Grains also hold the chemical equilibrium of formation versus dissociation inside disks, as their surfaces are the places where most $\mathrm{H_{2}}$ and $\mathrm{H_{2}O}$ are formed (e.g. 2002ApJ...575L..29C; 2004ApJ...604.. [^2] C; 2014ApJ...786..135A). Moreover, dust surfaces can adsorb gas-phase charge carriers (ions and free electrons) rather efficiently, letting these carriers deposit their free charge and get neutralized as they get desorbed (e.g. 2006A&A...445..205I; see also Bai+Goodman2009). This mechanism maintains a low but non-zero fraction of ionization in the pertinent zones in PPDs, enabling the non-ideal magnetohydrodynamic (MHD) effects to drive disk winds and accretion flows efficiently (Bai+Goodman2009; 2011ApJ...739...51B; 2017ApJ...845...75B; 2019ApJ...874...90W). Volatile coatings on dust grain surfaces, for example water, regulate grain sintering and thereby shape the dust size distribution throughout PPD evolution (2016ApJ...821...82O; 2017ApJ...845...31H). Such modulated distributions could influence their interactions with charged particles and lead to the formation of sub-structures in PPDs by alternating the magnetic diffusivities, the key parameters determining the interactions between the magnetic fields and the gas (e.g. 2019ApJ...885...36H; 2021ApJ...913..133H). More generally, the dynamic events related to dust growth, including coagulation, bouncing, breaking, and shattering of grains, are sensitively dependent on the sticking energy between surfaces (1993ApJ...407..806C; 1997ApJ...480..647D). Such sticking energy should also be susceptible to the conditions on the grain surfaces. Discussions regarding these events are extensive under various circumstances, including theoretical analyses (e.g. 2004ApJ...616..895Y; 2009MNRAS.394.1061H; 2009A&A...502..845O; 2014MNRAS.437.1636H), and even experimental studies (2000SSRv...92..265B).
@@
As the crucial characterization of adsorption processes, the energy of adsorption largely determines the status of volatile species and dust grains themselves in dust-rich astrophysical systems. It is widely assumed, based on observations and experiments, that the astrophysical dust grains can be broadly categorized into silicate grains and carbonaceous grains, whose mass fractions are in the same order of magnitude (see e.g. 1994A&A...292..641J; 1995A&A...300..503D; 1998A&A...332..291J; 2000A&A...364..282F; 2010ARA&A..48...21H; see also the data compilations in DraineBook).
@@
Over the past few decades, the adsorption condition of key species has been measured mainly by experimental studies. These experimental results should, nevertheless, be clarified and specified for their applicability and compatibility. For example, the reported adsorption energy of water on graphite is measured to be $E_{\mathrm{ad}}/k_{\mathrm{B}}\simeq 4800~{\rm K}$, almost identical to the adsorption on silicates (1990Icar...87..188S; 2004MNRAS.354.1141V; 2005JChPh.122d4713B; 2006A&A...445..205I; 2022ESC.....6..597M). These values neveretheless seem at odds with the quantum chemistry studies concluding that graphite is only mildly hydrophilic by adsorbing molecules with the van der Waals (vdW hereafter) interactions (see also PhysRevB.84.033402; hamada2012adsorption; brandenburg2019physisorption), while the silicate surfaces could adsorb water mainly by covalent or coordination bonds and hydrogen bonds. Additionally, experimental studies have focused on the single-species adsorption conditions only, while the adsorbates in PPDs could be a mixture of multiple molecular species. The impact of the surface coverage conditions of grains deserve further and more thorough discussions with modern theoretical and computational methods.
@@
In this paper, we will thoroughly study the problems related to the adsorption of molecules via computational chemistry. We will focus on the adsorption of water ($\mathrm{H_{2}O}$) and carbon monoxide ($\mathrm{CO}$), the most abundant molecules in the PPDs and molecular clouds (other than $\mathrm{H_{2}}$ and $\mathrm{He}$), onto the representative dust grain surface models, emphasizing the physical pictures of the results and compare them with experiments. This paper is structured as follows. §2 describes the ab-initio calculation methods of adsorption on carbonaceous and silicate grains, assuming single and multiple molecule adsorption scenarios. Using these data, §3 evaluates the surface coating conditions of both carbonaceous and silicate dust grains. §5 summarizes the paper, and extends the scope of the results to the gas-phase molecule condensation and sublimation processes onto grain surfaces, the potential impact on planet formation mechanisms.
@@
2 Ab-initio adsorption energy calculations
@@
The condensation process of key molecules onto astrophysical grain surfaces is calculated in this work by examining adsorption conditions using advanced computational chemistry techniques. The ab-initio density functional theory (DFT) package VASP (PhysRevB.47.558; PhysRevB.54.11169) is adopted, utilizing projector augmented wave (PAW) method calculations with PAW potentials for the atoms involved. To ensure accurate estimations of van der Waals (vdW) interactions, the combination of r2SCAN PhysRevLett.115.036402; furness2020accurate and rVV10 PhysRevB.87.041108 functionals is employed for electron exchange-correlation (XC) interactions. This combination, which involves generalized gradient approximations (GGA) and extensions to higher-order derivatives (meta-GGA), is confirmed by various tests and benchmarks (PhysRevX.6.041005) to handle vdW interactions under the DFT framework accurately in a wide range of situations, especially physisorptions. For instance, the single-molecule physisorption of water onto graphene, which is a highly relevant case to this work, is found to yield an adsorption energy $|\Delta\epsilon_{\mathrm{ad}}|\simeq 0.096~{\rm eV}$ when using r2SCAN+rVV10, a result that is remarkably close to fiducial quantum Monte Carlo (QMC) calculations (PhysRevB.84.033402; brandenburg2019physisorption). Within the PAW paradigm, a set of plane-wave bases is utilized to expand the electronic wave functions, with a cutoff energy of $600~{\rm eV}$.
@@
$\Delta\epsilon_{\mathrm{ad}}=\epsilon_{X}+\epsilon_{\mathrm{sub}}-\epsilon_{\mathrm{tot}}\ .$ (1)
@@
2.1 Adsorption onto Carbonaceous Surfaces
Carbonaceous dust grains are considered to exist in graphitic-like states, including graphites, oxidized graphites, graphene, and even polycyclic aromatic hydrocarbons (PAHs) at the smallest sizes $\lesssim 0.01~{\rm\mu m}$ (see also DraineBook), or as amorphous carbon. In terms of their astrophysical properties, these materials are observed to behave similarly, with the surface layer of carbon atoms dominating due to the weak vdW coupling between graphitic layers.
2.1.1 Single-molecule Adsorption onto Graphene
@@
Beginning with water molecules, following brandenburg2019physisorption, the molecules are positioned above the center of a carbon hexagon, with orientations classified as “0-leg”, “1-leg”, and “2-leg” (illustrated in Figure 2). For each orientation, a series of $\Delta E_{\rm abs}$ values is obtained at different displacements $d$ from the substrate surface, measured from the plane of carbon nuclei to the center-of-mass of the adsorbate molecule. It should be noted that during this step, the ionic structure of the system is fixed to manually iterate through configuration space. This setup ensured that the adsorbate configuration—particularly its orientation and distance to the substrates—remained unchanged, allowing correct energy curves to be obtained. The resulting energy curves for these configurations are presented in Figure 2. The difference between our approach (the r2SCAN + rVV10 method) and the QMC method for the 1-leg case (PhysRevB.84.033402) is $\lesssim 0.05~{\rm eV}$, well within the range of variations among different XC methods (hamada2012adsorption). Similar calculations for $\mathrm{H_{2}}$ and $\mathrm{CO}$ are performed following the same paradigm, with molecular orientation notations illustrated in Figure 2. For completeness and verification purposes, the adsorption of helium is also examined, yielding an adsorption energy of $|\Delta\epsilon_{\mathrm{ad}}|\lesssim 0.07~{\rm eV}$.
@@
2.1.2 Adsorption onto amorphous carbon surfaces
The realistic composition of interstellar and protostellar carbonaceous grains is considered to differ considerably from graphites or graphene. According to experimental studies 1998A&A...332..291J, these grains are more likely to be amorphous carbon, which in various cases may be doped with oxygen and hydrogen at approximately $\sim 1~\%$ of atomic concentration. Hydrogen atoms are expected to be located predominantly at grain surfaces due to their inability to achieve coordination numbers greater than unity. Oxygen atoms may exist in both the bulk substrate and near the surface; however, numerical experiments indicate that surface oxygen atoms are always saturated by capturing hydrogen atoms because of intense dipole moments. In practice, while the cases with doped atoms are also calculated, only hydrogen and carbon atoms are elaborated and explored in details. The construction of amorphous carbon substrates is explained in Appendix B.2.
@@ 2.1.2 Adsorption onto amorphous carbon surfaces
At each site, ionic structures are relaxed to locate the energy minimum before the adsorbate is moved sufficiently far from the substrate to evaluate $\Delta\epsilon_{\mathrm{ad}}$ using equation (1). Single-molecule adsorption energies in Table 1 are obtained by averaging $\Delta\epsilon_{\mathrm{ad}}$ over the five sampling sites. It is noted that replacing approximately $\sim 2~\%$ of carbon atoms with oxygen yielded nearly identical results, as oxygen atoms are rarely exposed directly (due to hydrogen surface coverage) and thus contributed minimally to interactions with adsorbates. Through deliberate exposure of oxygen and nitrogen atoms at the surface (e.g., Figure 3), the adsorption energy of $\mathrm{H_{2}O}$ via hydrogen bonds is measured to be $0.35~{\rm eV}$ (above oxygen) and $0.29~{\rm eV}$ (above nitrogen), respectively. The adsorption energy of $\mathrm{CO}$ near doping atoms exhibited no significant variations.
@@
When substrate surface coverage fractions become non-negligible, interactions between adsorbates may play a crucial role in modifying their kinetic behaviors. The vdW interactions and hydrogen bonds between adjacent molecules increase the barriers for desorption and hopping, with intensity dependent on both the numbers and types of neighbors. These effects are quantified through numerical experiments based on amorphous carbon substrates. The interaction energy is measured through the following procedures: (1) the adsorbate and its neighbors are placed above the substrate, and ionic configurations are evolved with DFT until stress relaxation, after which the total energy of the relaxed system is evaluated; (2) the adsorbate of interest is moved sufficiently far from the substrate ($\Delta z\gtrsim 10$$\text{\AA }$), and configurations of all remaining atoms are relaxed to reevaluate the total energy.
@@
Energy differences between these steps yielded the “neighbor-enhanced” total binding energy. For tests involving adsorbates with one neighbor, adsorbate pairs are positioned at five different sites (with centers of mass offsets following equation 2) to obtain average enhanced binding energies. The overall binding energy (including adsorption) is observed to increase sub-linearly, as vdW forces exhibit saturation behavior and hydrogen bonds require specific molecular orientations. Tests are also conducted to determine the numbers of different neighbors that maximize absolute binding energy values, beyond which binding energy decreases due to repulsive forces.
2.2 Adsorption onto Silicates
@@
In comparison to carbonaceous grains, silicate compositions are recognized to be considerably more complex. To maintain model tractability while preserving physical relevance, iron-poor enstatite $\mathrm{MgSiO_{3}}$ (Pnma space group using the Materials Project data, see also MatProj) is selected as a proxy for magnesium silicate substrates. This choice is motivated by the abundance of magnesium, which among alkali and alkaline earth metals in the ISM and PPDs, typically matches silicon abundance levels. To emulate astronomical silicate formation conditions, hydrogen atoms are added to exposed oxygen atoms on model silicate surfaces, following procedures analogous to those employed for amorphous carbon. The details of silicate substrate model construction are explained in Appendix B.3.
@@
The specific atomic composition of silicate substrates beneath adsorbate molecules is found to critically determine adsorption properties. Initial investigations positioned molecules with varying orientations at different heights above different atomic sites. After identifying configurations with approximately minimal energy, molecules are placed accordingly on the substrate, and ionic structures are evolved to achieve relaxation, eliminating total stress to the desired accuracy ($\lesssim 10^{-2}~{\rm eV}~\text{\AA }^{-1}$) and thereby locating energy extrema.
@@
Configurations converged at energy extrema are illustrated in Figure 4, with corresponding adsorption energies presented in Table 2. The adsorption energy values for $\mathrm{H_{2}O}$ and $\mathrm{CO}$ are generally observed to be one order of magnitude greater than those on carbonaceous surfaces, clearly indicating chemisorption processes. This enhancement is attributed to the coordination numbers of metal centers (Si atoms on $(001)^{-}$ and Mg atoms elsewhere) being significantly less than the saturation value of 6 in bulk $\mathrm{MgSiO_{3}}$, thereby permitting coordination bond formation and elevating adsorption energies to approximately $\sim 1~{\rm eV}$. A notable exception is identified for $\mathrm{H_{2}O}$ adsorption on the (100) surface, where exceptionally high adsorption energy resulted in molecular dissociation.
@@
Note. — Similar to Table 1, but the neighbor conditions are not included as the molecules are chemisorbed onto silicate surfaces (see also §2.2).
@@
3 Kinetic Monte Carlo Simulations of Grain Surfaces
Surface coverage conditions of dust grains play a fundamental role in regulating coagulation processes through their influence on sticking behavior and surface energy dynamics. To systematically investigate these coverage conditions, we developed a specialized kinetic Monte Carlo (KMC) simulation code, optimized for GPU acceleration to handle the computational demands of these complex simulations [^1] https://github.com/wll745881210/kmc_ads. whose general procedures are explained in Appendix C. It is also noted that the zero-temperature results obtained in §2 need thermodynamic corrections before being applied in the KMC simulations (see Appendix A).
3.1 Setups of KMC simulations

### v3

@@ I Introduction
The adsorption of volatile chemical species onto the surfaces of dust grains is of broad and critical interest to researchers studying the interstellar medium (ISM), star formation, and planet formation. For example, in protoplanetary disks (PPDs hereafter), dusts may play a central role in controlling the thermochemical conditions. Grains also hold the chemical equilibrium of formation versus dissociation inside disks, as their surfaces are the places where most $\mathrm{H_{2}}$ and $\mathrm{H_{2}O}$ are formed (e.g. Cazaux and Tielens, 2002, 2004; Ádámkovics et al., 2014). Moreover, dust surfaces can adsorb gas-phase charge carriers (ions and free electrons) rather efficiently, letting these carriers deposit their free charge and get neutralized as they get desorbed (e.g. Ilgner and Nelson 2006; see also Bai and Goodman 2009). This mechanism maintains a low but non-zero fraction of ionization in the pertinent zones in PPDs, enabling the non-ideal magnetohydrodynamic (MHD) effects to drive disk winds and accretion flows efficiently (Bai and Goodman, 2009; Bai, 2011, 2017; Wang et al., 2019). Volatile coatings on dust grain surfaces, for example water, regulate grain sintering and thereby shape the dust size distribution throughout PPD evolution (Okuzumi et al., 2016; Hasegawa et al., 2017). Such modulated distributions could influence their interactions with charged particles and lead to the formation of sub-structures in PPDs by alternating the magnetic diffusivities, the key parameters determining the interactions between the magnetic fields and the gas (e.g. Hu et al., 2019, 2021). More generally, the dynamic events related to dust growth, including coagulation, bouncing, breaking, and shattering of grains, are sensitively dependent on the sticking energy between surfaces (Chokshi et al., 1993; Dominik and Tielens, 1997). Such sticking energy should also be susceptible to the conditions on the grain surfaces. Discussions regarding these events are extensive under various circumstances, including theoretical analyses (e.g. Yan et al., 2004; Hirashita and Yan, 2009; Ormel et al., 2009; Hirashita and Voshchinnikov, 2014), and even experimental studies (Blum, 2000).
@@
As the crucial characterization of adsorption processes, the energy of adsorption largely determines the status of volatile species and dust grains themselves in dust-rich astrophysical systems. It is widely assumed, based on observations and experiments, that the astrophysical dust grains can be broadly categorized into silicate grains and carbonaceous grains, whose mass fractions are in the same order of magnitude (see e.g. Jaeger et al. 1994; Dorschner et al. 1995; Jager et al. 1998; Fabian et al. 2000; Henning 2010; see also the data compilations in Draine 2011).
@@
Over the past few decades, the adsorption condition of key species has been measured mainly by experimental studies. These experimental results should, nevertheless, be clarified and specified for their applicability and compatibility. For example, the reported adsorption energy of water on graphite is measured to be $E_{\mathrm{ad}}/k_{\mathrm{B}}\simeq 4800~{\rm K}$, almost identical to the adsorption on silicates (Sandford and Allamandola, 1990; Viti et al., 2004; Bolina et al., 2005; Ilgner and Nelson, 2006; Minissale et al., 2022). These values neveretheless seem at odds with the quantum chemistry studies concluding that graphite is only mildly hydrophilic by adsorbing molecules with the van der Waals (vdW hereafter) interactions (see also Ma et al., 2011; Hamada, 2012; Brandenburg et al., 2019), while the silicate surfaces could adsorb water mainly by covalent or coordination bonds and hydrogen bonds. Additionally, experimental studies have focused on the single-species adsorption conditions only, while the adsorbates in PPDs could be a mixture of multiple molecular species. The impact of the surface coverage conditions of grains deserve further and more thorough discussions with modern theoretical and computational methods.
@@
In this paper, we will thoroughly study the problems related to the adsorption of molecules via computational chemistry. We will focus on the adsorption of water ($\mathrm{H_{2}O}$) and carbon monoxide ($\mathrm{CO}$), the most abundant molecules in the PPDs and molecular clouds (other than $\mathrm{H_{2}}$ and $\mathrm{He}$), onto the representative dust grain surface models, emphasizing the physical pictures of the results and compare them with experiments. This paper is structured as follows. §II describes the ab-initio calculation methods of adsorption on carbonaceous and silicate grains, assuming single and multiple molecule adsorption scenarios. Using these data, §III evaluates the surface coating conditions of both carbonaceous and silicate dust grains. §V summarizes the paper, and extends the scope of the results to the gas-phase molecule condensation and sublimation processes onto grain surfaces, the potential impact on planet formation mechanisms.
@@
II Ab-initio adsorption energy calculations
@@
The condensation process of key molecules onto astrophysical grain surfaces is calculated in this work by examining adsorption conditions using advanced computational chemistry techniques. The ab-initio density functional theory (DFT) package VASP (Kresse and Hafner 1993; Kresse and Furthmüller 1996) is adopted, utilizing projector augmented wave (PAW) method calculations with PAW potentials for the atoms involved. To ensure accurate estimations of van der Waals (vdW) interactions, the combination of r2SCAN Sun et al. (2015); Furness et al. (2020) and rVV10 Sabatini et al. (2013) functionals is employed for electron exchange-correlation (XC) interactions. This combination, which involves generalized gradient approximations (GGA) and extensions to higher-order derivatives (meta-GGA), is confirmed by various tests and benchmarks (Peng et al., 2016) to handle vdW interactions under the DFT framework accurately in a wide range of situations, especially physisorptions. For instance, the single-molecule physisorption of water onto graphene, which is a highly relevant case to this work, is found to yield an adsorption energy $|\Delta\epsilon_{\mathrm{ad}}|\simeq 0.096~{\rm eV}$ when using r2SCAN+rVV10, a result that is remarkably close to fiducial quantum Monte Carlo (QMC) calculations (Ma et al. 2011; Brandenburg et al. 2019). Within the PAW paradigm, a set of plane-wave bases is utilized to expand the electronic wave functions, with a cutoff energy of $600~{\rm eV}$.
@@
$\Delta\epsilon_{\mathrm{ad}}=\epsilon_{\mathrm{tot}}-(\epsilon_{X}+\epsilon_{\mathrm{sub}})\ .$ (1)
@@
II.1 Adsorption onto Carbonaceous Surfaces
Carbonaceous dust grains are considered to exist in graphitic-like states, including graphites, oxidized graphites, graphene, and even polycyclic aromatic hydrocarbons (PAHs) at the smallest sizes $\lesssim 0.01~{\rm\mu m}$ (see also Draine 2011), or as amorphous carbon. In terms of their astrophysical properties, these materials are observed to behave similarly, with the surface layer of carbon atoms dominating due to the weak vdW coupling between graphitic layers.
II.1.1 Single-molecule Adsorption onto Graphene
@@
Beginning with water molecules, following Brandenburg et al. (2019), the molecules are positioned above the center of a carbon hexagon, with orientations classified as “0-leg”, “1-leg”, and “2-leg” (illustrated in Figure 2). For each orientation, a series of $\Delta E_{\rm abs}$ values is obtained at different displacements $d$ from the substrate surface, measured from the plane of carbon nuclei to the center-of-mass of the adsorbate molecule. It should be noted that during this step, the ionic structure of the system is fixed to manually iterate through configuration space. This setup ensured that the adsorbate configuration—particularly its orientation and distance to the substrates—remained unchanged, allowing correct energy curves to be obtained. The resulting energy curves for these configurations are presented in Figure 2. The difference between our approach (the r2SCAN + rVV10 method) and the QMC method for the 1-leg case (Ma et al., 2011) is $\lesssim 0.05~{\rm eV}$, well within the range of variations among different XC methods (Hamada, 2012). Similar calculations for $\mathrm{H_{2}}$ and $\mathrm{CO}$ are performed following the same paradigm, with molecular orientation notations illustrated in Figure 2. For completeness and verification purposes, the adsorption of helium is also examined, yielding an adsorption energy of $|\Delta\epsilon_{\mathrm{ad}}|\lesssim 0.07~{\rm eV}$.
@@
II.1.2 Adsorption onto amorphous carbon surfaces
The realistic composition of interstellar and protostellar carbonaceous grains is considered to differ considerably from graphites or graphene. According to experimental studies Jager et al. (1998), these grains are more likely to be amorphous carbon, which in various cases may be doped with oxygen and hydrogen at approximately $\sim 1~\%$ of atomic concentration. Hydrogen atoms are expected to be located predominantly at grain surfaces due to their inability to achieve coordination numbers greater than unity. Oxygen atoms may exist in both the bulk substrate and near the surface; however, numerical experiments indicate that surface oxygen atoms are always saturated by capturing hydrogen atoms because of intense dipole moments. In practice, while the cases with doped atoms are also calculated, only hydrogen and carbon atoms are elaborated and explored in details. The construction of amorphous carbon substrates is explained in Appendix B.2.
@@ II.1.2 Adsorption onto amorphous carbon surfaces
At each site, ionic structures are relaxed to locate the energy minimum before the adsorbate is moved sufficiently far from the substrate to evaluate $\Delta\epsilon_{\mathrm{ad}}$ using equation (1). Single-molecule adsorption energies in Table 1 are obtained by averaging $\Delta\epsilon_{\mathrm{ad}}$ over the five representative sampling sites. It is acknowledged that the inherent stochasticity and structural heterogeneity of amorphous carbon surfaces could be better represented by a larger ensemble of sampling points characterizing such disordered media. Nonetheless, an exhaustive sampling of the near-infinite configuration space of local adsorption sites is computationally prohibitive. The amorphous carbon substrate model generated for this study (Appendix B.2) also exhibits some degree of structural homogeneity within the supercell that facilitates representative sampling of the dominant binding environments. In fact, supplementary DFT calculations have also been performed on a wider set of prospective sites and alternative amorphous carbon realizations (not shown in this paper), whose consequent adsorption energies consistently fall within the range established by the five primary sampling points (except a few extreme cases, where the adsorbates eventually reside in the sites whose $|\Delta\epsilon_{\rm ad}|$ still stay in the consistent range after structure relaxation).
@@
In It is noted that replacing approximately $\sim 2~\%$ of carbon atoms with oxygen yielded nearly identical results, as oxygen atoms are rarely exposed directly (due to hydrogen surface coverage) and thus contributed minimally to interactions with adsorbates. Through deliberate exposure of oxygen and nitrogen atoms at the surface (e.g., Figure 3), the adsorption energy of $\mathrm{H_{2}O}$ via hydrogen bonds is measured to be $0.35~{\rm eV}$ (above oxygen) and $0.29~{\rm eV}$ (above nitrogen), respectively. The adsorption energy of $\mathrm{CO}$ near doping atoms exhibited no significant variations.
@@
When substrate surface coverage fractions become non-negligible, interactions between adsorbates may play a crucial role in modifying their kinetic behaviors. The vdW interactions and hydrogen bonds between adjacent molecules increase the barriers for desorption and hopping, with intensity dependent on both the numbers and types of neighbors. These effects are quantified through numerical experiments based on amorphous carbon substrates. The interaction energy is measured through the following procedures:
  1. 1.
    Initial site pre-optimization: The “primary” adsorbate of interest is first placed and relaxed at a previously descreibed single-molecule adsorption site via full geometry optimization.
  2. 2.
    Initial neighbor positioning: Neighboring molecules are introduced at radial distances approximating the characteristic intermolecular separations found in their respective bulk solid phases (e.g., the hydrogen-bonding network of amorphous solid water when the primary molecule and its neighbors are all $\mathrm{H_{2}O}$). The azimuthal angles of these neighbors are selected so that the neighbors are close to their likely single-molecule adsorption site if possible (e.g., the nearest available hexagonal centers or “hollow” sites over the amorphous surface), or in random otherwise (while the separations between neighbors are still guaranteed).
  3. 3.
    Constrained adsorbate relaxation: DFT structural optimization is conducted where the degrees of freedom of all adsorbed molecules are relaxed to a local energy minimum. During this step, the coordinates of the substrate atoms are held fixed to maintain the pre-equilibrated surface morphology.
  4. 4.
    Adsorption energy measurement: Adsorbate of interest is moved sufficiently far from the substrate ($\Delta z\gtrsim 10$$\text{\AA }$), and configurations of all remaining atoms are relaxed to re-evaluate the total energy.
@@
Energy differences measured through these steps yielded the “neighbor-enhanced” total binding energy. For tests involving adsorbates with one neighbor, adsorbate pairs are positioned at five different sites (with centers of mass offsets following equation 2) to obtain average enhanced binding energies. The overall binding energy (including adsorption) is observed to increase sub-linearly, as vdW forces exhibit saturation behavior and hydrogen bonds require specific molecular orientations. Tests are also conducted to determine the numbers of different neighbors that maximize absolute binding energy values, beyond which binding energy decreases due to repulsive forces. Similar to the single-molecule cases, depleting the possibilities over the heterogeneous amorphous carbon surfaces is virtually not practical. Nonetheless, because of the qualitative homogeneity of the model amorphous surface, it is expected that such procedures could still reasonably sample the typical situations that are relevant in the following calculations.
II.2 Adsorption onto Silicates
@@
In comparison to carbonaceous grains, silicate compositions are recognized to be considerably more complex. To maintain model tractability while preserving physical relevance, iron-poor enstatite $\mathrm{MgSiO_{3}}$ (Pnma space group using the Materials Project data, see also Jain et al. 2013) is selected as a proxy for magnesium silicate substrates. This choice is motivated by the abundance of magnesium, which among alkali and alkaline earth metals in the ISM and PPDs, typically matches silicon abundance levels.The details of silicate substrate model construction are explained in Appendix B.3.
Admittedly, the model crystalline Pnma enstatite is unable to fully represent the amorphous enstatite, a commonly encountered phase of astrosilicates also in PPDs (see Henning, 2010, and references therein). Practical representation of disordered systems in the ab-initio DFT frameworks, nevertheless, typically involves significant computational trade-offs. While crystalline substrates can be accurately modeled with relatively compact supercells ($\sim 10^{2}$ atoms), a physically realistic amorphous structure requires substantially larger, disordered models (typically $\sim 10^{3}$ atoms) to minimize periodic artifacts, making systematic potential energy surface mapping with advanced metaGGA and van der Waals functionals (e.g., r2SCAN+rVV10 adopted in this work) prohibitively expensive.
In this study, crystalline enstatite is adopted as a generic proxy for its amorphous counterpart, becaus (1) both phases share very similar stoichiometry, providing a chemically representative building block for magnesium-rich silicates; and (2) surface reactivity is primarily governed by local coordination environments rather than long-range order, with adsorption dominated by undercoordinated Mg and Si sites whose chemical interactions remain consistent across phases. To better approximate the morphological and chemical heterogeneity of a disordered surface within these computational constraints, calculations were performed across multiple distinct facets including (100), (010), and $(001)^{\pm}$, which expose varying ratios of Mg and Si atoms in diverse coordination states, effectively bracketing the range of local environments encountered on a realistic silicate surface.
@@
The specific atomic composition of silicate substrates beneath adsorbate molecules is found to determine adsorption properties. Initial investigations positioned molecules with varying orientations at different heights above different atomic sites. After identifying configurations with approximately minimal energy, molecules are placed accordingly on the substrate, and ionic structures are evolved to achieve relaxation, eliminating total stress to the desired accuracy ($\lesssim 10^{-2}~{\rm eV}~\text{\AA }^{-1}$) and thereby locating energy extrema. In astrophysical conditions relevant to PPDs, it is possible that the silicate surfaces could be covered by chemisorbed hydrogen. Additional tests were performed on surfaces partially covered with these hydrogen (not shown in this paper), which preferentially saturates exposed under-coordinated oxygen atoms while leaving some Mg and Si coordination centers accessible. These tests indicate that hydrogen coverage does not significantly alter the adsorption energies of $\mathrm{H_{2}O}$ and CO on enstatite surfaces (with changes limited to $\lesssim 0.2~{\rm eV}$), which is of secondary importance compared to the total chemisorption energy of $|\Delta\epsilon_{\rm ad}|\gtrsim 1~{\rm eV}$. This is consistent with expectation, as non-closest-neighbor conditions do not strongly affect covalent coordinate bonds.
@@
Configurations converged at energy extrema are illustrated in Figure 4, with corresponding adsorption energies presented in Table 2. The adsorption energy values for $\mathrm{H_{2}O}$ and $\mathrm{CO}$ are generally observed to be one order of magnitude greater than those on carbonaceous surfaces, clearly indicating chemisorption processes. This enhancement is attributed to the coordination numbers of metal centers (Si atoms on $(001)^{-}$ and Mg atoms elsewhere) being significantly less than the saturation value of 6 in bulk $\mathrm{MgSiO_{3}}$, thereby permitting coordination bond formation and elevating adsorption energies to approximately $\sim 1~{\rm eV}$. A notable exception is identified for $\mathrm{H_{2}O}$ adsorption on the (100) surface where the molecule adsorption eventuall leads to dissociation. On this facet, exposed magnesium atoms exhibit a preferred directionality for coordination bond formation that lies nearly tangential to the surface plane, positioning the chemisorbed water molecule close to neighboring under-coordinated silicon atoms (with coordination numbers of $\sim 4$, versus 6 in the bulk). These Si sites could accept and stabilize the transferred proton, collectively lowering the activation barrier for O–H bond cleavage. In comparison, the (010), $(001)^{+}$, and $(001)^{-}$ surfaces favor coordination geometries that extend more perpendicularly from the surface, keeping sufficient distances of the water molecule from under-coordinated atoms, while exposing more fully coordinated Si–O–Si bridging configurations that are less susceptible to protonation.
@@
Note. — Similar to Table 1, but the neighbor conditions are not included as the molecules are chemisorbed onto silicate surfaces (see also §II.2).
@@
III Kinetic Monte Carlo Simulations of Grain Surfaces
Surface coverage conditions of dust grains play a fundamental role in regulating coagulation processes through their influence on sticking behavior and surface energy dynamics. To systematically investigate these coverage conditions, we developed a specialized kinetic Monte Carlo (KMC) simulation code, optimized for GPU acceleration to handle the computational demands of these complex simulations [^1] https://github.com/wll745881210/kmc_ads. whose general procedures are explained in Appendix C. The zero-temperature results obtained in §II need thermodynamic corrections before being applied in the KMC simulations (see Appendix A). It is also noticed that the KMC approach itself inherently sacrifices the accuracy of ab-initio calculations by discretizing events and transition rates on a square lattice, rather than employing rigorous off-lattice simulations that directly utilize consistent DFT calculations. In this work, the transition rates are also approximate values derived from limited sampling of substrate conditions, based on adsorption and binding energies obtained from a limited-size amorphous carbon model or from crystalline enstatite as a proxy for amorphous silicate. Nevertheless, the purpose of conducting KMC simulations is to balance this trade-off, to obtain improved statistical sampling over larger systems at the cost of single-process accuracy. By using test results based on a model amorphous carbon surface with reasonable homogeneity, and by bracketing amorphous silicate surface environments with multiple crystalline facets, the KMC results are expected to retain reasonable semi-quantitative reliability.
III.1 Setups of KMC simulations